# Adaptive Coded Federated Learning: Privacy Preservation and Straggler Mitigation

### v1

@@ III-B The second stage: training iterations
$I_{t}^{\left(i\right)}=\begin{cases}1,&\text{if device }i\text{ is not a straggler},\\ 0,&\text{if device }i\text{ is a straggler}.\end{cases}$ (5)
@@
and then transmits ${\mathbf{G}}_{t}^{\left(i\right)}$ to the central server. At the same time, the central server computes a gradient based on the global coded dataset to obtain:
@@
where $\alpha_{t}\in[0,1]$ represents the adaptive aggregation weights. Here, in the proposed method, we introduce an adaptive policy for adjusting the values of $\alpha_{t}$ over iterations to attain the optimal performance in terms of privacy and learning, which will be specified after conducting performance analysis in the next section. With the global gradient provided in (9), the central server updates the model parameter matrix as
@@
and broadcasts ${{\mathbf{W}}_{t+1}}$ to all devices, which ends the $t$-th iteration. In (10), $\eta_{t}$ denotes the learning rate.
@@
$\psi_{2}^{(T)}$ $=\phi(Tod)N,$ (13)
where $\phi$ denotes the number of bits required for the transmission of a real number. In many practical applications, the FL process involves a very large number of iterations with $T\gg d$, which indicates $\psi_{2}^{(T)}\gg\psi_{1}^{(T)}$. Hence, it holds that ${\psi^{(T)}}\approx\psi_{2}^{(T)}$. In other words, compared with conventional FL only involving training iterations [8], the extra communication overhead induced by the first stage of the proposed method is negligible.
@@ IV Performance Analysis and The Adaptive Policy
In this section, we first analyze the performance of the proposed method with arbitrary aggregation weights from two aspects: privacy performance and learning performance. Based on that, to attain the optimal performance in terms of privacy and learning, the adaptive policy of determining the value of $\alpha_{t}$ would be specified accordingly. First, let us provide some useful assumptions, which have been widely and reasonably adopted in the literature.
@@ IV-A Privacy performance of ACFL
To evaluate the privacy performance of the proposed method, the metric of MI-DP is adopted, which is stronger than the conventional differential privacy metric [31, 32]. Let us provide the definition of MI-DP as follows [25]:
@@
The matrix $q\left({\mathbf{Z}}\right)={\mathbf{\tilde{Z}}}\in{\mathbb{R}^{{a_{2}}\times{b_{2}}}}$ encoded from ${\mathbf{Z}}\in{\mathbb{R}^{{a_{1}}\times{b_{1}}}}$ with randomized function $q(\cdot)$ satisfies $\epsilon$-MI-DP if
$\sup_{j,k,p({\mathbf{Z}})}I(Z_{j,k};q({\mathbf{Z}})\|\mathbf{Z}^{-j,k})\leq\epsilon,$ (16)
where $Z_{j,k}$ denotes the $(j,k)$-th element in ${\mathbf{Z}}$, $\mathbf{Z}^{-j,k}$ is the set of all elements in ${\mathbf{Z}}$ excluding $Z_{j,k}$, and $p({\mathbf{Z}})$ denotes the distribution of ${\mathbf{Z}}$. A decreasing value of $\epsilon$ indicates enhanced privacy performance.
@@
$\mathop{\sup}\limits_{j,k,p\left({{\mathbf{X}},{\mathbf{Y}}}\right)}I\left({\left.{X_{j,k};{\mathbf{H}}_{X},{\mathbf{H}}_{Y}}\right|{{\mathbf{X}}^{-j,k}},{\mathbf{Y}}}\right)$
$=$ $\mathop{\sup}\limits_{p\left({{\mathbf{X}},{\mathbf{Y}}}\right)}I\left({\left.{X_{1,1};{\mathbf{H}}_{X},{\mathbf{H}}_{Y}}\right|{{\mathbf{X}}^{-1,1}},{\mathbf{Y}}}\right),$ (18)
it is sufficient to show
$\mathop{\sup}\limits_{p\left({{\mathbf{X}},{\mathbf{Y}}}\right)}I\left({\left.{X_{1,1};{\mathbf{H}}_{X},{\mathbf{H}}_{Y}}\right|{{\mathbf{X}}^{-1,1}},{\mathbf{Y}}}\right)\leq\epsilon,$ (19)
$\mathop{\sup}\limits_{p\left({{\mathbf{X}},{\mathbf{Y}}}\right)}I\left({\left.{X_{1,1};{\mathbf{H}}_{X},{\mathbf{H}}_{Y}}\right|{{\mathbf{X}}^{-1,1}},{\mathbf{Y}}}\right)$
$=$ $\mathop{\sup}\limits_{p\left({{\mathbf{X}},{\mathbf{Y}}}\right)}I\left({\left.{X_{1,1};{{\mathbf{X}}^{T}}{\mathbf{X}}+{\mathbf{N}}_{1},{{\mathbf{X}}^{T}}{\mathbf{Y}}+{\mathbf{N}}_{2}}\right|{{\mathbf{X}}^{-1,1}},{\mathbf{Y}}}\right)$
$=$ $\mathop{\sup}\limits_{p\left({{\mathbf{X}},{\mathbf{Y}}}\right)}h\left({\left.{{{\mathbf{X}}^{T}}{\mathbf{X}}+{\mathbf{N}}_{1},{{\mathbf{X}}^{T}}{\mathbf{Y}}+{\mathbf{N}}_{2}}\right|{{\mathbf{X}}^{-1,1}},{\mathbf{Y}}}\right)$
$=$ $\mathop{\sup}\limits_{p\left({{\mathbf{X}},{\mathbf{Y}}}\right)}h\left({\left.{{{\mathbf{X}}^{T}}{\mathbf{X}}+{\mathbf{N}}_{1},{{\mathbf{X}}^{T}}{\mathbf{Y}}+{\mathbf{N}}_{2}}\right|{{\mathbf{X}}^{-1,1}},{\mathbf{Y}}}\right)$
$-h\left({{\mathbf{N}}_{1},{\mathbf{N}}_{2}}\right),$ (20)
where $h\left({\cdot}\right)$ denotes the entropy, the first equality is derived by substituting (2) and (3) into (IV-A) and the second equality holds according to the definition of mutual information. In (IV-A), for the first term, we have (IV-A) based on the independence between the additive noise and the raw data.
$\mathop{\sup}\limits_{p\left({{\mathbf{X}},{\mathbf{Y}}}\right)}h\left({\left.{{{\mathbf{X}}^{T}}{\mathbf{X}}+{\mathbf{N}}_{1},{{\mathbf{X}}^{T}}{\mathbf{Y}}+{\mathbf{N}}_{2}}\right|{{\mathbf{X}}^{-1,1}},{\mathbf{Y}}}\right)$
$=$ $\mathop{\sup}\limits_{p\left({{\mathbf{X}},{\mathbf{Y}}}\right)}h\left({\left.{\left\{{{X_{1,j}}{X_{1,1}}+{{\left({{N_{1}}}\right)}_{j,1}},j\neq 1}\right\},\left\{{{X_{1,k}}{X_{1,1}}+{{\left({{N_{1}}}\right)}_{1,k}},k\neq 1}\right\},X_{1,1}^{2}+{{\left({{N_{1}}}\right)}_{1,1}},\left\{{{X_{1,1}}{Y_{1,k}}+{{\left({{N_{2}}}\right)}_{1,k}},\forall k}\right\}}\right|{{\mathbf{X}}^{-1,1}},{\mathbf{Y}}}\right)$
$\leq$ $\mathop{\sup}\limits_{p\left({\mathbf{X}}\right)}\sum\limits_{j\neq 1}{h\left({\left.{{X_{1,j}}{X_{1,1}}+{{\left({{N_{1}}}\right)}_{j,1}}}\right|{{\mathbf{X}}^{-1,1}}}\right)}+\mathop{\sup}\limits_{p\left({\mathbf{X}}\right)}\sum\limits_{k\neq 1}{h\left({\left.{{X_{1,k}}{X_{1,1}}+{{\left({{N_{1}}}\right)}_{1,k}}}\right|{{\mathbf{X}}^{-1,1}}}\right)}+\mathop{\sup}\limits_{p\left({{X_{1,1}}}\right)}h\left({X_{1,1}^{2}+{{\left({{N_{1}}}\right)}_{1,1}}}\right)$
$+\mathop{\sup}\limits_{p\left({\mathbf{Y}}\right)}\sum\limits_{k}{h\left({\left.{{X_{1,1}}{Y_{1,k}}+{{\left({{N_{2}}}\right)}_{1,k}}}\right|{\mathbf{Y}}}\right)}+h\left({\left\{{{{\left({{{N}}_{1}}\right)}_{j,k}},j\neq 1,k\neq 1}\right\}}\right)+h\left({\left\{{{{\left({{{N}}_{2}}\right)}_{j,k}},j\neq 1,\forall k}\right\}}\right)$
$\leq$ $\mathop{\sup}\limits_{p\left({\mathbf{X}}\right)}\sum\limits_{j\neq 1}{h\left({{X_{1,j}}{X_{1,1}}+{{\left({{N_{1}}}\right)}_{j,1}}}\right)}+\mathop{\sup}\limits_{p\left({\mathbf{X}}\right)}\sum\limits_{k\neq 1}{h\left({{X_{1,k}}{X_{1,1}}+{{\left({{N_{1}}}\right)}_{1,k}}}\right)}+\mathop{\sup}\limits_{p\left({{X_{1,1}}}\right)}h\left({X_{1,1}^{2}+{{\left({{N_{1}}}\right)}_{1,1}}}\right)$
$+\mathop{\sup}\limits_{p\left({\mathbf{Y}}\right)}\sum\limits_{k}{h\left({{X_{1,1}}{Y_{1,k}}+{{\left({{N_{2}}}\right)}_{1,k}}}\right)}+h\left({\left\{{{{\left({{{N}}_{1}}\right)}_{j,k}},j\neq 1,k\neq 1}\right\}}\right)+h\left({\left\{{{{\left({{{N}}_{2}}\right)}_{j,k}},j\neq 1,\forall k}\right\}}\right).$ (21)
Substituting (IV-A) into (IV-A) yields
$\mathop{\sup}\limits_{p\left({{\mathbf{X}},{\mathbf{Y}}}\right)}I\left({\left.{X_{1,1};{\mathbf{H}}_{X},{\mathbf{H}}_{Y}}\right|{{\mathbf{X}}^{-1,1}},{\mathbf{Y}}}\right)$
$\leq$ $\mathop{\sup}\limits_{p\left({\mathbf{X}}\right)}\sum\limits_{j\neq 1}{h\left({{X_{1,j}}{X_{1,1}}+{{\left({{N_{1}}}\right)}_{j,1}}}\right)}$
$+\mathop{\sup}\limits_{p\left({\mathbf{X}}\right)}\sum\limits_{k\neq 1}{h\left({{X_{1,k}}{X_{1,1}}+{{\left({{N_{1}}}\right)}_{1,k}}}\right)}$
$+\mathop{\sup}\limits_{p\left({{X_{1,1}}}\right)}h\left({X_{1,1}^{2}+{{\left({{N_{1}}}\right)}_{1,1}}}\right)$
$+\mathop{\sup}\limits_{p\left({\mathbf{Y}}\right)}\sum\limits_{k}{h\left({{X_{1,1}}{Y_{1,k}}+{{\left({{N_{2}}}\right)}_{1,k}}}\right)}$
$-h\left({\left\{{{{\left({{{N}}_{1}}\right)}_{j,k}},j=1,k\neq 1}\right\}}\right)$
$-h\left({\left\{{{{\left({{{N}}_{1}}\right)}_{j,k}},j\neq 1,k=1}\right\}}\right)$
based on Assumption 1 and the maximum entropy bound [33]. From (IV-A), we can easily derive Theorem 1. ∎
@@ IV-B Learning performance of ACFL
${\mathbf{G}}_{t}^{True}=\sum\limits_{i=1}^{N}{{\mathbf{G}}_{t}^{\left(i\right)}},$ (24)
@@
$\mathbb{E}\left({\left.{{\mathbf{G}}_{t}^{S}}\right|{{\mathbf{W}}_{t}}}\right)=\mathbb{E}\left({{{{\mathbf{\tilde{H}}}}_{X}}}\right){{\mathbf{W}}_{t}}-\mathbb{E}\left({{{{\mathbf{\tilde{H}}}}_{Y}}}\right)={\mathbf{G}}_{t}^{True}.$ (25)
@@
$={\alpha_{t}}{\mathbf{G}}_{t}^{True}+\left({1-{\alpha_{t}}}\right)\sum\limits_{i=1}^{N}{{\mathbf{G}}_{t}^{\left(i\right)}}={\mathbf{G}}_{t}^{True},$ (26)
@@
We can derive that
based on (2), (3), (4), (8), (9) and the fact that the additive noise is independent from the straggler behaviour of the devices. For the first term in (IV-B), we have
×$\times\mathbb{E}\left[{\left({{\alpha_{t}}+\frac{{1-{\alpha_{t}}}}{{1-p}}I_{t}^{\left({{i_{1}}}\right)}}\right)\left({{\alpha_{t}}+\frac{{1-{\alpha_{t}}}}{{1-p}}I_{t}^{\left({{i_{2}}}\right)}}\right)}\right]$
and the second inequality holds due to Assumption 2. For the second term in (IV-B), we have
$=\alpha_{t}^{2}Nd\sigma_{1}^{2}\left\|{{{\mathbf{W}}_{t}}}\right\|_{F}^{2}$
$\leq\alpha_{t}^{2}Nd\sigma_{1}^{2}{C^{2}},$ (31)
based on the independence of the additive noise on different local datasets and Assumption 3. For the third term in (IV-B), we have
Substituting (IV-B), (IV-B) and (32) into (IV-B) yields Lemma 2. ∎
@@
[…]
@@ IV-C The adaptive policy
which aims to expedite the convergence of the learning process under certain privacy level with fixed variances of additive noise. Based on the optimization problem defined in (35), we detail the adaptive policy for selecting the value of $\alpha_{t}$ and analyze the trade off between privacy and learning performance under this policy in Theorem 3 as follows.
@@
$+N{\beta^{2}}\frac{1}{{1-p}}+N{\beta^{2}}\left({N-1}\right),$ (38)
@@
Noting that $\frac{{pN{\beta^{2}}\frac{1}{{1-p}}}}{{\left({pN{\beta^{2}}\frac{1}{{1-p}}+Nd\sigma_{1}^{2}{C^{2}}+N\sigma_{2}^{2}od}\right)}}<1$ always holds, which indicates that
@@
According to Theorem 3, the learning performance of ACFL improves as the variances of the additive noise decrease, whereas the privacy performance benefits from an increase in these variances. This highlights the intrinsic trade off between privacy and learning performance. Practically, the variances of the additive noise can be meticulously calibrated to achieve a desired trade off, aligning with specific practical requirements.
To verify that ACFL with the adaptive policy achieves a better performance in terms of privacy and learning, Fig. 1 illustrates the relationship between the upper bound derived in (33) and $\epsilon$ specified in (17) with non-adaptive and fixed values of aggregation weights as well as adaptive aggregation weights determined by the designed adaptive policy, where we set $d=100,\sigma_{1}^{2}=\sigma^{2}_{2},o=10,p=0.1,N=5,\beta=10,c=1,T=1000$ and $\lambda=1$. Here, a lower value of the upper bound implies superior learning performance, whereas a lower $\epsilon$ value signifies enhanced privacy performance. The results clearly show that under equivalent privacy levels, the upper bound in (33) reaches its minimum with the adaptive policy. This outcome indicates that ACFL, when implemented with the adaptive policy, attains the optimal performance in terms of privacy and learning.
Figure 1: Trade off between the learning and privacy performance. (a) The relationship between the upper bound in (33) and $\epsilon$ for ACFL under non-adaptive aggregation weights and adaptive aggregation weights. (b) The relationship between the upper bound in (33) and $\epsilon$ for ACFL under adaptive aggregation weights determined by the adaptive policy.
In Theorem 3, we develop the adaptive policy for setting the aggregation weights in ACFL, which relies on the bounds specified in Assumptions 2 and 3. Notably, obtaining the values of $C$ and $\beta$ beforehand can be challenging in certain scenarios. In such cases, a practical implementation of ACFL involves estimating $C$ and $\beta$ over the iterations by the central server. Specifically, $C$ and $\beta$ are approximated by $\hat{C}$ and $\hat{\beta}$, respectively, with their expressions provided as follows:
@@
With (43), the implementation of ACFL is presented in Algorithm 1 and shown as Fig. 2.
Figure 2: The implementation of ACFL.

### v2

@@ III-B The second stage: training iterations
[…]
@@
and then transmits ${\mathbf{G}}_{t}^{\left(i\right)}$ to the central server. From (7), we can clearly see that $\mathbf{X}^{(i)T}\mathbf{X}^{(i)}$ and $\mathbf{X}^{(i)T}\mathbf{Y}^{(i)}$ contain information equivalent to that in the raw datasets for computing gradients in the considered problem, which implies that it is reasonable to project the local datasets as described in (2) and (3) to generate the coded datasets. At the same time, the central server computes a gradient based on the global coded dataset to obtain:
@@
where $\alpha_{t}\in[0,1]$ represents the adaptive aggregation weight. For the proposed method, we will introduce an adaptive policy for adjusting the value of $\alpha_{t}$ over iterations to attain the optimal performance in terms of privacy and learning after conducting performance analysis in the next section. With the global gradient provided in (9), the central server updates the model parameter matrix as
@@
and broadcasts ${\mathbf{W}}_{t+1}$} to all devices, which ends the $t$-th iteration. In (10), $\eta_{t}$ denotes the learning rate. The implementation of the proposed method is shown as Fig. 3.
Figure 3: The implementation of ACFL.
@@
$\psi_{2}^{(T)}$ $=\phi od\sum\limits_{i=1}^{N}{\sum\limits_{t=0}^{T}{I_{t}^{\left(i\right)}}},$ (13)
where $\phi$ denotes the number of bits required for the transmission of a real number. In many practical applications, the FL process involves a large number of iterations with $T\gg\frac{{d+o}}{{o\left({1-p}\right)}}$, which can be verified under our simulation settings in Section V, indicating $\psi_{2}^{(T)}\approx\phi odN(1-p)\left({T+1}\right)\gg\psi_{1}^{(T)}$. Hence, it holds that ${\psi^{(T)}}\approx\psi_{2}^{(T)}$. Based on this, the additional communication overhead introduced by the first stage of the proposed method is negligible compared to the overall communication cost of the training process.
Note that for conventional FL involving only training iterations [10], the communication overhead arises from uploading gradients from the devices to the central server in each iteration. In this context, the per-iteration communication overhead of conventional FL and ACFL is equivalent. Given that the additional communication overhead introduced by the first stage of ACFL is negligible compared to the communication cost incurred during training iterations, both conventional FL and ACFL incur approximately the same overall communication burden after the same number of iterations.
@@ IV Performance Analysis and The Adaptive Policy
In this section, we first analyze the performance of the proposed method with arbitrary aggregation weights from two aspects: privacy performance and learning performance. Based on that, to attain the optimal performance in terms of privacy and learning, the adaptive policy of determining the value of $\alpha_{t}$ would be specified accordingly. Note that, in this paper, we focus on analyzing the privacy performance during the sharing of local coded datasets with the server, rather than the transmission of local gradients during training iterations, similar as [31, 29]. This focus is motivated by the observation that, compared with the typical FL framework, the additional privacy concerns introduced by the proposed method arise only in the first stage, where local coded datasets are shared with the server. In contrast, the privacy concerns associated with sharing local gradients exist universally across all FL scenarios. Based on that, it is sufficient to analyze the additional privacy concerns in the ACFL framework to evaluate the advantages and limitations of the proposed method. Additionally, numerous works in the FL literature address privacy concerns related to sharing local gradients with the server [38]. These techniques can be seamlessly incorporated into the proposed method to enhance privacy during the second stage. However, this is beyond the primary scope of this work. In light of this, in this paper, when we refer to the privacy performance of ACFL, we specifically mean the privacy performance of ACFL in sharing local coded datasets with the server.
First, let us provide some useful assumptions, which have been widely and reasonably adopted in the literature.
@@ IV-A Privacy performance of ACFL
To evaluate the privacy performance of the proposed method in sharing local coded datasets with the server, the metric of MI-DP is adopted, which is stronger than the conventional differential privacy metric [39, 40]. Let us provide the definition of MI-DP as follows [32]:
@@
The coded dataset $q\left({\mathbf{Z}}\right)={\mathbf{\tilde{Z}}}\in{\mathbb{R}^{{a_{2}}\times{b_{2}}}}$ obtained from the original dataset ${\mathbf{Z}}\in{\mathbb{R}^{{a_{1}}\times{b_{1}}}}$ with encoding operation $q(\cdot)$ satisfies $\epsilon$-MI-DP if
$\sup_{j,k,\Gamma({\mathbf{Z}})}I(Z_{j,k};q({\mathbf{Z}})\|\mathbf{Z}^{-j,k})\leq\epsilon,$ (16)
where $Z_{j,k}$ denotes the $(j,k)$-th element in $\mathbf{Z}$}, $\mathbf{Z}^{-j,k}$ is the set of all elements in $\mathbf{Z}$} excluding $Z_{j,k}$, $\Gamma({\mathbf{Z}})$ denotes the distribution of $\mathbf{Z}$}, and $I(\cdot||\cdot)$ is the conditional mutual information.
In this definition, $\epsilon$ is an upper bound of the maximum privacy leakage of each element in the original dataset $\mathbf{Z}$} given the coded dataset $q\left({\mathbf{Z}}\right)$. A decreasing value of $\epsilon$ indicates enhanced privacy performance.
@@
Please see Appendix A. ∎
@@
Remark 3.
In the typical FL setting, gathering the training data of all devices at the server is challenging due to two main reasons: (1) sharing raw training data directly with the server raises privacy concerns, especially when the data is sensitive, and (2) the communication overhead of transmitting local datasets to the server is significant.
In contrast, the proposed ACFL framework addresses both issues by allowing devices to share local coded datasets with the server. First, privacy concerns are mitigated by transforming the original datasets with additive noise before transmission. This ensures that the privacy leakage caused by transmitting data can be well controlled, as shown in Theorem 1 and Remark 2. Second, the communication overhead is significantly reduced by transmitting local coded datasets instead of raw datasets, as explained in Section III-A.
@@ IV-B Learning performance of ACFL
${\mathbf{G}}_{t}^{\text{True}}=\sum\limits_{i=1}^{N}{{\mathbf{G}}_{t}^{\left(i\right)}},$ (19)
@@
$\mathbb{E}\left({\left.{{\mathbf{G}}_{t}^{S}}\right|{{\mathbf{W}}_{t}}}\right)=\mathbb{E}\left({{{{\mathbf{\tilde{H}}}}_{X}}}\right){{\mathbf{W}}_{t}}-\mathbb{E}\left({{{{\mathbf{\tilde{H}}}}_{Y}}}\right)={\mathbf{G}}_{t}^{\text{True}}.$ (20)
@@
$={\alpha_{t}}{\mathbf{G}}_{t}^{\text{True}}+\left({1-{\alpha_{t}}}\right)\sum\limits_{i=1}^{N}{{\mathbf{G}}_{t}^{\left(i\right)}}={\mathbf{G}}_{t}^{\text{True}},$ (21)
@@
Please see Appendix B. ∎
@@
Note that the optimal solution for the considered linear regression problem is given as
$\mathbf{W}^{*}=\left(\sum_{i=1}^{N}\mathbf{X}^{(i)T}\mathbf{X}^{(i)}\right)^{-1}\left(\sum_{i=1}^{N}\mathbf{X}^{(i)T}\mathbf{Y}^{(i)}\right),$ (25)
which relies solely on the matrices ${{\mathbf{X}}^{\left(i\right)T}}{{\mathbf{X}}^{\left(i\right)}}$ and ${{\mathbf{X}}^{\left(i\right)T}}{{\mathbf{Y}}^{\left(i\right)}}$.
Based on (25), the server can obtain an approximate version of the exact optimal solution using the received local coded datasets $\mathbf{H}_{X}^{(i)}$ and $\mathbf{H}_{Y}^{(i)}$ in the first stage, as defined in (2) and (3). In other words, the central server receives a noisy version of the matrices ${{\mathbf{X}}^{\left(i\right)T}}{{\mathbf{X}}^{\left(i\right)}}$ and ${{\mathbf{X}}^{\left(i\right)T}}{{\mathbf{Y}}^{\left(i\right)}}$ in the first stage and can determine an approximate solution immediately. However, the error in this approximate solution can become arbitrarily large as the strength of the additive noise increases, which corresponds to an enhanced privacy performance of ACFL, as shown in Remark 2. When there is a stringent requirement for privacy performance, this approximate solution becomes inapplicable. Besides, in many applications, it is essential to acquire the exact solution rather than an approximate one. For instance, in medical data analysis, where the linear regression models are used to predict diseases, an exact solution ensures accurate diagnostics. Small errors in the solution could lead to incorrect diagnoses.
In contrast, as observed from Theorem 2, the proposed method converges to the optimal solution exactly, providing an exact and reliable solution at any privacy level. Based on that, employing the proposed method to obtain an exact optimal solution is more practical than directly deriving an approximate solution from the local coded datasets.
@@
Note that the condition $\sum_{i=1}^{N}\mathbf{X}^{(i)T}\mathbf{X}^{(i)}\geq\lambda\mathbf{I}$ with $\lambda>0$ in Theorem 2 naturally holds for the considered problem formulated in Section II. This is because $\mathbf{X}^{(i)}\in\mathbb{R}^{m_{i}\times d}$ is full column rank, and $\mathbf{X}^{(i)T}\mathbf{X}^{(i)}$ is positive definite. In this case, $\lambda>0$ can be any value that satisfies $0<\lambda<\sum_{i=1}^{N}\text{eig}_{\min}\left[\mathbf{X}^{(i)T}\mathbf{X}^{(i)}\right]$, where $\text{eig}_{\min}(\cdot)$ denotes the smallest eigenvalue of a matrix.
@@ IV-C The adaptive policy
which aims to expedite the convergence of the learning process under certain privacy level with fixed variances of additive noise. Based on the optimization problem defined in (26), we detail the adaptive policy for selecting the value of $\alpha_{t}$ and analyze the privacy and learning performance under this policy in Theorem 3 as follows.
@@
$+N{\beta^{2}}\frac{1}{{1-p}}+N{\beta^{2}}\left({N-1}\right).$ (29)
@@
We note that $\frac{{pN{\beta^{2}}\frac{1}{{1-p}}}}{{{pN{\beta^{2}}\frac{1}{{1-p}}+Nd\sigma_{1}^{2}{C^{2}}+N\sigma_{2}^{2}od}}}<1$ always holds, which indicates that
@@
From Theorem 3, we observe that the aggregation weight corresponding to the gradient computed by the server based on the global coded dataset decreases as the strength of the additive noise increases. This aligns with our intuition that when stronger noise is added to the local coded datasets, the gradient derived from the global coded dataset becomes less accurate and contains more noise, making it less reliable for updating the global model. Additionally, when the straggler probability $p$ increases, devices are more likely to become stragglers, leading to a greater loss of information. With more missing information from the stragglers, the local gradients from the non-stragglers deviate further from the true gradient, making it more important to leverage the information provided by the server to compensate for the missing contributions. As a result, the aggregation weight of the gradient computed by the server based on the global coded dataset increases.
@@
According to Theorem 1 and Theorem 3, the learning performance of ACFL improves as the variances of the additive noise decrease, whereas the privacy performance benefits from an increase in these variances. This highlights the intrinsic trade-off between privacy and learning performance. Practically, the variances of the additive noise can be meticulously calibrated to achieve a desired trade-off, aligning with specific practical requirements.
To verify that ACFL with the adaptive policy achieves a better performance in terms of privacy and learning, Fig. 4 illustrates the relationship between the upper bound derived in (23) and $\epsilon$ specified in (17) with non-adaptive and fixed values of aggregation weights as well as adaptive aggregation weights determined by the designed adaptive policy, where we set $d=100,\sigma_{1}^{2}=\sigma^{2}_{2},o=10,p=0.1,N=5,\beta=10,c=1,T=1000$ and $\lambda=1$. Here, a lower value of the upper bound implies superior learning performance, whereas a lower $\epsilon$ value signifies enhanced privacy performance. The results clearly show that under the same privacy level, the upper bound in (23) reaches its minimum with the adaptive policy. This outcome indicates that ACFL, when implemented with the adaptive policy, attains the optimal performance in terms of privacy and learning.
Figure 4: The trade-off between the learning and privacy performance is shown by depicting the upper bound in (23) as a function of $\epsilon$ for ACFL under non-adaptive aggregation weights and adaptive aggregation weights.
In Theorem 3, we develop the adaptive policy for setting the aggregation weights in ACFL, which relies on the bounds specified in Assumptions 2 and 3. Notably, obtaining the values of $C$ and $\beta$ beforehand can be challenging in certain scenarios. In such cases, a practical implementation of ACFL is estimating $C$ and $\beta$ over the iterations by the central server. Specifically, $C$ and $\beta$ are approximated by $\hat{C}$ and $\hat{\beta}$, respectively, with their expressions provided as follows:
@@
With (34), the implementation of ACFL is presented in Algorithm 1. While the use of estimated bounds in (33) may lead to a slight degradation from theoretically optimal performance, our simulation results in Section V will demonstrate that the proposed method remains convergent and exhibits stable behavior across all tested scenarios. This indicates that the discrepancy between the theoretical analysis and practical implementation has a negligible impact on the stability and final performance of the algorithm. Moreover, the method consistently outperforms existing baselines, suggesting that the use of estimated bounds is sufficiently robust in practice.
From (33) and (34), it can be seen that the computational load associated with computing the aggregation weights adaptively is $\mathcal{O}\left(odN(1-p)\right)$. Based on this, the server-side weight computation time increases with the size of the model parameter matrix. However, this additional computational overhead does not necessarily indicate a delay in the overall algorithm, for the following reasons. First, the server is typically equipped with sufficient computational resources to handle the additional load efficiently. As a result, the extra computation time incurred at the server is negligible when compared to the total computational cost of other components of the algorithm during training. Second, the server can begin processing as soon as it starts receiving messages from the fastest non-straggler device. This enables the server-side computation to proceed in parallel with ongoing device-side computation and communication. In other words, the server does not need to wait until all non-straggler devices have completed their transmissions before beginning its computations. This parallelism implies that the additional computation time introduced by computing the aggregation weights is negligible when compared to the baseline case in which fixed weights are employed by the server.